OpenStudy (anonymous):

Can someone tell me if I am correct....Asking to find all #'s for which the expression is undefined 6 over 8w+1 I think this is a trick and the expression is defined for all real #'s, right?

OpenStudy (chaise):

\[\frac{6}{8w+1}\] When the fraction is undefined: \[8w+1=0\] \[8w=-1\] \[w=-1/8\] When w = 1/8 the fraction is undefined as the denominator is equal to zero.

OpenStudy (anonymous):

-1/8 or positive?

OpenStudy (chaise):

Negative. Sorry. When you multiply (-1/8)x8 you get -1. Add one to this and you get 0 (and undefined) Sorry for any confusion.